sides of rectangle are in 5:2 if the width of the rectangle is 5cm, then find the area of rectangle
Answer by mananth(16946)

let length be l
5/2 = l/5
l=25/2
l= 12.5
...
Area = 12.5 * 5= 62.5 cm^2
...
m.ananth@hotmail.ca